OpenStudy (anonymous):

Multiply and simplify: (2x-1)(3x+4)

hero (hero):

(2x - 1)(3x + 4) = 2x(3x + 4) - 1(3x + 4) = 6x^2 + 8x - 3x - 4 = 6x^2 + 5x - 4

OpenStudy (anonymous):

where u get the 5x from?

OpenStudy (anonymous):

never mind I see it

OpenStudy (anonymous):

thanks